Samples of size 8 are selected from a manufacturing process. The mean of the sample ranges is 0.6. What is the estimate of the standard deviation of the population? (Round your answer to 3 decimal places.)
Answer #1

We have given,

Mean =np......................Binomial probability distribution mean formula.

0.6=8*p = > p = 0.075

Therefore,

X\sim Binomial(n=8,p=0.075)

Standard deviation =\sqrt{npq}=\sqrt{8*0.075*(1-0.075)}=0.745

Standard deviation =0.745